eing able to read and analyze financial statements is an important skill for managers and investors. In this assignment you will calculate a number of measures derived from a simple income statement and balance sheet. Income Statement Item Value Sales $5,234,198 Cost of Goods Sold $3,659,135 Gross Profit $1,575,063
Selling Expense $227,295 General & Administrative $826,340 Total Operating Expense $1,053,635
Operating Profit $521,428 Less Income Tax (40%) $208,571 Net Profit $312,857
Dividends Paid (100k shares) $100,000
Continued on next page...
Student or Team Name:
Balance Sheet Account Value Cash $165,852 Accounts Receivable $589,367 Inventory $1,765,129
Plant & Equipment $5,500,000 Less Depreciation $1,100,000 Net Plant & Equipment $4,400,000
Total Assets $6,920,348
Account Value Accounts Payable $496,231 Long-Term Debt $2,950,000 Total Liabilities $3,446,231
Common Stock $2,000,000 Retained Earnings $1,474,117 Total Equity $3,474,117
Total Liability & Equity $6,920,348
Use the financial statements above to calculate the following measures: Measure Formula Value 1. Current Ratio current assets ÷ current liabilities 2. Debt to Equity Ratio total debt ÷ total equity 3. Gross Margin gross profit ÷ sales 4. Return on Sales (ROS) net profit ÷ sales 5. Return on Equity (ROE) net profit ÷ total equity 6. Earnings per Share (EPS) net profit ÷ shares outstandin
Answer #1

1. Current Ratio current = assets ÷ current liabilities = (165852 + 589367 + 1765129 ) / 496231 = 5.08

2. Debt to Equity Ratio = total debt ÷ total equity = 3446231 / 3474117 = 0.992

3. Gross Margin = gross profit ÷ sales = 1,575,063 / 5,234,198 = 30.09%

4. Return on Sales (ROS) = net profit ÷ sales = 312857 / 5,234,198 = 5.98%

5. Return on Equity (ROE) = net profit ÷ total equity = 312857 / 3474117 = 9%

6. Earnings per Share (EPS) = net profit ÷ shares outstanding = 312857 / 100000 = 3.13%